# The dark side of the universe may be more harmonic than we thought

### v2

@@ I Introduction
In the PAge-like unified dark fluid (PUDF) model that was proposed in Wang et al. [49], the unified dark component is assumed to be a fluid with a smooth background evolution parameterized by the PAge approximation [56, 57]. The PAge approximation is based on two assumptions, that the dark component(s) behave like dust at high-redshift, and that the dimensionless combination $Ht$, where $H$ is the Hubble parameter and $t$ is the age of the universe, is a slowly varying smooth function of $t$. The minimal PUDF contains seven cosmological parameters, with the standard $\Omega_{c}h^{2}$ (CDM density) replaced by the PAge parameters $p_{\rm age}$ ($\sim$ age of the universe) and $\eta$ (deviation from Einstein-de Sitter universe). By modifying the Boltzmann code CLASS [58], Wang et al. [49] computed the linear perturbations in PUDF and found that PUDF can give predictions similar to those of $\Lambda$CDM. Further analysis of Bayesian evidence shows that $\Lambda$CDM is favored over PUDF by the current cosmological data including CMB, BAO, SNe, and cosmic chronometers (CC) [49].
The results found in Wang et al. [49], however, lack a clear physical interpretation. It is unclear to what extent PUDF can mimic $\Lambda$CDM at the background and linear-perturbation levels. Neither do we know what key difference between PUDF and $\Lambda$CDM has led to the slightly different $\chi^{2}$ fits to the data. Similar problems exist for the earlier work Davari et al. [55] with a polynomial-based parameterization. This work then aims to improve the theoretical understanding of the similarities and nuances between PUDF and $\Lambda$CDM, and to come up with some quantitative predictions that can be used to test the numerical accuracy of the Boltzmann code. While the theoretical exploration is done in Section II, we revisit the Bayesian parameter inference and update some of the results in Section III. Section IV summarizes and concludes.
@@
II Theoretical Comparison between PUDF and $\Lambda$CDM
II.1 PUDF basics
@@
$H^{2}(z)$ $=$ $H_{0}^{2}\left[\Omega_{\gamma}+\sum_{i=1}^{3}\Omega_{\nu,i}\frac{\mathrm{I}_{\rho}\left(\frac{m_{\nu,i}}{(1+z)T_{\nu}}\right)}{\mathrm{I}_{\rho}\left(\frac{m_{\nu,i}}{T_{\nu}}\right)}\right](1+z)^{4}$ (1)
@@
$\mathrm{I}_{\rho}(\lambda)\equiv\frac{1}{2\pi^{2}}\int_{0}^{\infty}\frac{x^{2}\sqrt{x^{2}+\lambda^{2}}}{e^{x}+1}~{}\mathrm{d}x.$ (2)
@@
$H_{\mathrm{PAge}}$ $=$ $H_{0}\sqrt{1-\Omega_{\nu}-\Omega_{\gamma}}$ (3)
@@
$1+z$ $=$ $\left(\frac{p_{\rm age}}{\beta}\right)^{2/3}$ (4)
@@
$\rho_{d}(z)=\frac{3}{8\pi G}H_{\mathrm{PAge}}^{2}-\rho_{b}(z),$ (5)
@@
$\rho_{b}(z)=\rho_{\rm crit}\Omega_{b}(1+z)^{3}\propto\Omega_{b}h^{2}(1+z)^{3}.$ (6)
@@
$w\equiv\frac{p_{d}}{\rho_{d}}=\frac{1+z}{3\rho_{d}}\frac{\mathrm{d}\rho_{d}}{\mathrm{d}z}-1.$ (8)
@@
$\dot{\delta}$ $=$ $-(1+w)\left(\theta+\frac{\dot{h}_{i}^{i}}{2}\right)-3\frac{\dot{a}}{a}\left(c_{\mathrm{s,eff}}^{2}-w\right)\delta$ (9)
@@
$\dot{\theta}$ $=$ $-\frac{\dot{a}}{a}\,\left(1-3c_{\mathrm{s,eff}}^{2}\right)\theta+\frac{c_{\mathrm{s,eff}}^{2}}{1+w}\,k^{2}\delta-k^{2}\sigma\,,$ (10)
@@
$c_{\mathrm{s,ad}}^{2}=\frac{\dot{P}}{\dot{\rho}}=w-\frac{\dot{w}}{3H\left(1+w\right)},$ (11)
@@
II.2 Matching the primary CMB
In the high-redshift limit where $\beta\sim H_{0}t\ll 1$, we may expand Eqs. (3-4) to the linear order of $\beta$ and obtain
$H_{\rm PAge}^{2}$$\approx$ $\frac{4H_{0}^{2}(1-\Omega_{\nu}-\Omega_{\gamma})}{9p_{\rm age}^{2}}e^{2+\eta-3p_{\rm age}}(1+z)^{3}$ (12)
@@ II.2 Matching the primary CMB
$\Omega_{c,\rm eff}=\frac{4(1-\Omega_{\nu}-\Omega_{\gamma})}{9p_{\rm age}^{2}}e^{2+\eta-3p_{\rm age}}-\Omega_{b},$ (13)
@@
$\rho_{d}\rvert_{\text{high }z}\approx\rho_{\rm crit}\Omega_{c,\rm eff}(1+z)^{3}.$ (14)
@@
$\left.\frac{4(1-\Omega_{\nu}-\Omega_{\gamma})}{9p_{\rm age}^{2}}e^{2+\eta-3p_{\rm age}}\right|_{\rm PUDF}=\left.\Omega_{m}\right|_{\rm\Lambda CDM},$ (15)
@@
$\left.\frac{4}{9p_{\rm age}^{2}}e^{2+\eta-3p_{\rm age}}\right|_{\rm PUDF}=\left.\Omega_{m}\right|_{\rm\Lambda CDM},$ (16)
@@
We use Eq. (15) to test the modified Boltzmann code CLASS in Wang et al. [49] and find an $O(10^{-3})$ relative difference between PUDF and $\Lambda$CDM primary CMB power spectra. Further investigation shows that this inconsistency is due to the usage of the subpackage HyRec, which contains a hard-coded $w_{0}w_{a}$ cosmology and therefore can be incompatible with modifications in CLASS. To fix this problem, we replace HyRec with the adapted version of RecFAST in CLASS, which reads cosmology from CLASS. The updated code agrees well with the theoretical expectation that once Eq. (15) is satisfied, the relative difference in primary CMB power spectra of PUDF and $\Lambda$CDM does not exceed $O(10^{-4})$. Figure 1 shows an example where PUDF is matched to the Planck 2018 bestfit $\Lambda$CDM model [1].
@@
Figure 1: Comparison of the primary CMB $TT$ and $EE$ power spectra of PUDF and $\Lambda$CDM when the matching condition (15) is applied. The lower panels give the relative difference ($\Delta C_{\ell}^{\mathrm{TT}}=C^{\mathrm{TT}}_{\ell,\mathrm{PUDF}}-C^{\mathrm{TT}}_{\ell,\Lambda\mathrm{CDM}},\,\Delta C_{\ell}^{\mathrm{EE}}=C^{\mathrm{EE}}_{\ell,\mathrm{PUDF}}-C^{\mathrm{EE}}_{\ell,\Lambda\mathrm{CDM}}$).
II.3 Matching late-universe observables
@@
For a given $\Omega_{m}\rvert_{\rm\Lambda CDM}$, Eq. (16) does not fix $p_{\rm age}$ and $\eta$. We may choose another constraint to match more observables between PUDF and $\Lambda$CDM. For instance, we may match the deceleration parameter $q_{0}=\frac{a\ddot{a}}{\dot{a}^{2}}$ in PUDF and $\Lambda$CDM. In the case of negligible $\Omega_{\nu}$ and $\Omega_{\gamma}$, the $q_{0}$ matching condition is
$\left.\frac{4(1-\eta)}{9p_{\rm age}^{2}}\right|_{\rm PUDF}=\left.\Omega_{m}\right|_{\rm\Lambda CDM}.$ (17)
@@
In the original work on PAge where only late universe observables were used, the primary-CMB matching condition (16) was not considered. Instead, the age of the universe in unit of $H_{0}$ was matched [56]. Ignoring the radiation and neutrinos, the age matching condition is
$\left.p_{\rm age}\right|_{\rm PUDF}=\left.\frac{2}{3\sqrt{1-\Omega_{m}}}\ln\frac{1+\sqrt{1-\Omega_{m}}}{\sqrt{\Omega_{m}}}\right|_{\rm\Lambda CDM}.$ (18)
@@
Figure 2: The primary-CMB matching condition (16), $q_{0}$ matching condition (17) and age matching condition (18) for $\Omega_{m}=0.28$ (left panel), $\Omega_{m}=0.315$ (middle panel) and $\Omega_{m}=0.35$ (right panel), respectively.
@@
In Figure 2 we plot the matching conditions for primary CMB, $q_{0}$ and age for a few representative $\Omega_{m}$ values. It is nontrivial to observe that the three conditions almost intersect at one point, where both early- and late-universe observables match well between PUDF and $\Lambda$CDM. It has been shown in Huang [56] that BAO and SN observables can be matched to percent-level accuracy between PAge and $\Lambda$CDM.
@@
While the background evolution is matched between PUDF and $\Lambda$CDM, the abundance and equation of state of the unified dark fluid in PUDF are very different from those of dark matter in $\Lambda$CDM. We may expect very different density perturbations of the dark components in the two models. However, density perturbations of the dark components are not directly observable. What can be observed are the density perturbations of baryonic matter and the bending of the light due to gravitational lensing, both of which track the gravitational potential $\phi$ if anisotropic stress can be ignored. The linear growth of $\phi$ in general depends on the total density perturbation $\delta\rho_{\rm tot}$, the total pressure perturbation $\delta p_{\rm tot}$, and the expansion history of the universe [59]. On sub-horizon scales where the gauge-dependence of $\delta\rho_{\rm tot}$ and $\delta p_{\rm tot}$ can be ignored, we may use the Poisson equation to eliminate the dependence on $\delta\rho_{\rm tot}$ [59]. Thus, in models such as PUDF and $\Lambda$CDM where the rest-frame pressure perturbations are assumed to be negligible, the evolution of $\phi$ on sub-horizon scales only depend on the expansion history of the universe. In other words, for background-matched PUDF and $\Lambda$CDM, the linear growth of gravitational potential is also approximately matched. This has been numerically verified in Wang et al. [49] where the baryon power spectrum in PUDF was shown to be similar to that in $\Lambda$CDM. In Figure 3 we show that the CMB lensing deflection power spectrum in PUDF and $\Lambda$CDM are similar, too.
@@
Figure 3: Comparison of CMB deflection power spectrum $C_{\ell}^{dd}$ of PUDF and $\Lambda$CDM when the primary-CMB matching condition (15) and the $q_{0}$ matching condition (17) are applied. The lower panel gives the relative difference($\Delta C_{\ell}^{\mathrm{dd}}=C^{\mathrm{dd}}_{\ell,\mathrm{PUDF}}-C^{\mathrm{dd}}_{\ell,\Lambda\mathrm{CDM}}$).
III Parameter Inference
@@
In this section we update the parameter inference for PUDF after fixing the minor error in the Boltzmann code. For a fair comparison we use the same combination of CMB + BAO + SN + CC that have been used in Wang et al. [49]. These datasets include Planck TTTEEE and lensing likelihoods [60, 61], the Pantheon+ compilation of Type Ia supernovae [62], and the recent Dark Energy Spectroscopic Instrument Data Release 1 (DESI DR1, Adame et al. [8]). More details can be found in Wang et al. [49].
@@
The results are listed in Table 2. Compared to the results in Wang et al. [49], the updated PUDF parameters all shift towards $\Lambda$CDM parameters. This is because in Wang et al. [49] the incorrect usage of HyRec leads to mismatched primary CMB and hence biased cosmological parameters. The updated results still favor $\Lambda$CDM model, but the relative Bayesian evidence derived with the MCEvidence code [63, 64] is updated to $\ln B_{\rm\Lambda CDM,PUDF}=3.75$, much less significant than the previous (incorrect) result $\ln B_{\rm\Lambda CDM,PUDF}=6.27$ in Wang et al. [49]. With a Bayesian evidence $\Delta\ln B\approx 3.75$, we may say that $\Lambda$CDM is only mildly preferred by the data and PUDF remains to be an interesting option for future investigation.
@@
data sets CMB+BAO+SN+CC CMB+BAO+CC CMB+SN+CC BBN+BAO+SN+CC
$\chi^{2}_{\rm PUDF}-\chi^{2}_{\rm\Lambda CDM}$ $6.5$ $0.58$ $1.1$ $-1.73$
IV Discussion and Conclusions
In this study we show that both the background expansion history and the linear perturbations in the visible sector of the universe can be tuned to be $\Lambda$CDM-like in the PAge-like unified dark fluid model. We derive matching conditions for primary CMB and late-universe observables. Using the primary-CMB matching condition to test the numerical accuracy of the PUDF Boltzmann code, we find a minor error in the code used in Wang et al. [49]. After fixing the numerical error we update the parameter inference and find that $\Lambda$CDM is mildly favored by the current CMB+BAO+SN+CC data, with a $\Delta\chi^{2}\approx 6.5$.

### v3

@@ 1 Introduction
In the PAge-like unified dark fluid (PUDF) model that was proposed in Wang:2024rus [49], the unified dark component is assumed to be a fluid with a smooth background evolution parameterized by the PAge approximation [56, 57]. The PAge approximation is based on two assumptions, that the dark component(s) behave like dust at high-redshift, and that the dimensionless combination $Ht$, where $H$ is the Hubble parameter and $t$ is the age of the universe, is a slowly varying smooth function of $t$. The minimal PUDF contains seven cosmological parameters, with the standard $\Omega_{c}h^{2}$ (CDM density) replaced by the PAge parameters $p_{\rm age}$ ($\sim$ age of the universe) and $\eta$ (deviation from Einstein-de Sitter universe). By modifying the Boltzmann code CLASS [58], Wang:2024rus [49] computed the linear perturbations in PUDF and found that PUDF can give predictions similar to those of $\Lambda$CDM. Further analysis of Bayesian evidence shows that $\Lambda$CDM is favored over PUDF by the current cosmological data including CMB, BAO, SNe, and cosmic chronometers (CC) [49].
The results found in Wang:2024rus [49], however, lack a clear physical interpretation. It is unclear to what extent PUDF can mimic $\Lambda$CDM at the background and linear-perturbation levels. Neither do we know what key difference between PUDF and $\Lambda$CDM has led to the slightly different $\chi^{2}$ fits to the data. Similar problems exist for the earlier work Davari2018 [55] with a polynomial-based parameterization. This work then aims to improve the theoretical understanding of the similarities and nuances between PUDF and $\Lambda$CDM, and to come up with some quantitative predictions that can be used to test the numerical accuracy of the Boltzmann code. While the theoretical exploration is done in Section 2, we revisit the Bayesian parameter inference and update some of the results in Section 3. Section 4 summarizes and concludes.
@@
2 Theoretical Comparison between PUDF and $\Lambda$CDM
2.1 PUDF basics
@@
$H^{2}(z)$ $=$ $H_{0}^{2}\left[\Omega_{\gamma}+\sum_{i=1}^{3}\Omega_{\nu,i}\frac{\mathrm{I}_{\rho}\left(\frac{m_{\nu,i}}{(1+z)T_{\nu}}\right)}{\mathrm{I}_{\rho}\left(\frac{m_{\nu,i}}{T_{\nu}}\right)}\right](1+z)^{4}$ (2.1)
@@
$\mathrm{I}_{\rho}(\lambda)\equiv\frac{1}{2\pi^{2}}\int_{0}^{\infty}\frac{x^{2}\sqrt{x^{2}+\lambda^{2}}}{e^{x}+1}~{}\mathrm{d}x.$ (2.2)
@@
$H_{\mathrm{PAge}}$ $=$ $H_{0}\sqrt{1-\Omega_{\nu}-\Omega_{\gamma}}$ (2.3)
@@
$1+z$ $=$ $\left(\frac{p_{\rm age}}{\beta}\right)^{2/3}$ (2.4)
@@
$\rho_{d}(z)=\frac{3}{8\pi G}H_{\mathrm{PAge}}^{2}-\rho_{b}(z),$ (2.5)
@@
$\rho_{b}(z)=\rho_{\rm crit}\Omega_{b}(1+z)^{3}\propto\Omega_{b}h^{2}(1+z)^{3}.$ (2.6)
@@
$w\equiv\frac{p_{d}}{\rho_{d}}=\frac{1+z}{3\rho_{d}}\frac{\mathrm{d}\rho_{d}}{\mathrm{d}z}-1.$ (2.8)
While Equations (2.3-2.8) may appear rather complicated, they stem from just two straightforward assumptions [56]: (1) the dark component behaves like dust at high redshifts $z\gg 1$, and (2) the dimensionless quantity $Ht$ can be approximated by a quadratic function of $t$. The Taylor expansion of $Ht$ is conceptually similar to expanding the dark energy equation of state $w(a)$, but the PUDF framework offers two key advantages over the $w_{0}w_{a}$ model: it is more economical, requiring one fewer parameter, and better physically motivated, as $Ht$ has been shown to vary smoothly and slowly across a wide range of well-studied cosmological models [56].
@@
$\dot{\delta}$ $=$ $-(1+w)\left(\theta+\frac{\dot{h}_{i}^{i}}{2}\right)-3\frac{\dot{a}}{a}\left(c_{\mathrm{s,eff}}^{2}-w\right)\delta$ (2.9)
@@
$\dot{\theta}$ $=$ $-\frac{\dot{a}}{a}\,\left(1-3c_{\mathrm{s,eff}}^{2}\right)\theta+\frac{c_{\mathrm{s,eff}}^{2}}{1+w}\,k^{2}\delta-k^{2}\sigma\,,$ (2.10)
@@
$c_{\mathrm{s,ad}}^{2}=\frac{\dot{P}}{\dot{\rho}}=w-\frac{\dot{w}}{3H\left(1+w\right)},$ (2.11)
@@
2.2 Matching the primary CMB
In the high-redshift limit where $\beta\sim H_{0}t\ll 1$, we may expand Eqs. (2.3-2.4) to the linear order of $\beta$ and obtain
$H_{\rm PAge}^{2}$$\approx$ $\frac{4H_{0}^{2}(1-\Omega_{\nu}-\Omega_{\gamma})}{9p_{\rm age}^{2}}e^{2+\eta-3p_{\rm age}}(1+z)^{3}$ (2.12)
@@ 2.2 Matching the primary CMB
$\Omega_{c,\rm eff}=\frac{4(1-\Omega_{\nu}-\Omega_{\gamma})}{9p_{\rm age}^{2}}e^{2+\eta-3p_{\rm age}}-\Omega_{b},$ (2.13)
@@
$\rho_{d}\rvert_{\text{high }z}\approx\rho_{\rm crit}\Omega_{c,\rm eff}(1+z)^{3}.$ (2.14)
@@
$\left.\frac{4(1-\Omega_{\nu}-\Omega_{\gamma})}{9p_{\rm age}^{2}}e^{2+\eta-3p_{\rm age}}\right|_{\rm PUDF}=\left.\Omega_{m}\right|_{\rm\Lambda CDM},$ (2.15)
@@
$\left.\frac{4}{9p_{\rm age}^{2}}e^{2+\eta-3p_{\rm age}}\right|_{\rm PUDF}=\left.\Omega_{m}\right|_{\rm\Lambda CDM},$ (2.16)
@@
We use Eq. (2.15) to test the modified Boltzmann code CLASS in Wang:2024rus [49] and find an $O(10^{-3})$ relative difference between PUDF and $\Lambda$CDM primary CMB power spectra. Further investigation shows that this inconsistency is due to the usage of the subpackage HyRec, which contains a hard-coded $w_{0}w_{a}$ cosmology and therefore can be incompatible with modifications in CLASS. To fix this problem, we replace HyRec with the adapted version of RecFAST in CLASS, which reads cosmology from CLASS. The updated code agrees well with the theoretical expectation that once Eq. (2.15) is satisfied, the relative difference in primary CMB power spectra of PUDF and $\Lambda$CDM does not exceed $O(10^{-4})$. Figure 1 shows an example where PUDF is matched to the Planck 2018 bestfit $\Lambda$CDM model [1].
@@
Figure 1: Comparison of the primary CMB $TT$ and $EE$ power spectra of PUDF and $\Lambda$CDM when the matching condition (2.15) is applied. The lower panels give the relative difference ($\Delta C_{\ell}^{\mathrm{TT}}=C^{\mathrm{TT}}_{\ell,\mathrm{PUDF}}-C^{\mathrm{TT}}_{\ell,\Lambda\mathrm{CDM}},\,\Delta C_{\ell}^{\mathrm{EE}}=C^{\mathrm{EE}}_{\ell,\mathrm{PUDF}}-C^{\mathrm{EE}}_{\ell,\Lambda\mathrm{CDM}}$).
2.3 Matching late-universe observables
@@
For a given $\Omega_{m}\rvert_{\rm\Lambda CDM}$, Eq. (2.16) does not fix $p_{\rm age}$ and $\eta$. We may choose another constraint to match more observables between PUDF and $\Lambda$CDM. For instance, we may match the deceleration parameter $q_{0}=\frac{a\ddot{a}}{\dot{a}^{2}}$ in PUDF and $\Lambda$CDM. In the case of negligible $\Omega_{\nu}$ and $\Omega_{\gamma}$, the $q_{0}$ matching condition is
$\left.\frac{4(1-\eta)}{9p_{\rm age}^{2}}\right|_{\rm PUDF}=\left.\Omega_{m}\right|_{\rm\Lambda CDM}.$ (2.17)
@@
In the original work on PAge where only late universe observables were used, the primary-CMB matching condition (2.16) was not considered. Instead, the age of the universe in unit of $H_{0}$ was matched [56]. Ignoring the radiation and neutrinos, the age matching condition is
$\left.p_{\rm age}\right|_{\rm PUDF}=\left.\frac{2}{3\sqrt{1-\Omega_{m}}}\ln\frac{1+\sqrt{1-\Omega_{m}}}{\sqrt{\Omega_{m}}}\right|_{\rm\Lambda CDM}.$ (2.18)
@@
Figure 2: The primary-CMB matching condition (2.16), $q_{0}$ matching condition (2.17) and age matching condition (2.18) for $\Omega_{m}=0.28$ (left panel), $\Omega_{m}=0.315$ (middle panel) and $\Omega_{m}=0.35$ (right panel), respectively.
@@
In Figure 2 we plot the matching conditions for primary CMB, $q_{0}$ and age for a few representative $\Omega_{m}$ values. It is nontrivial to observe that the three conditions almost intersect at one point, where both early- and late-universe observables match well between PUDF and $\Lambda$CDM. It has been shown in Huang:2020mub [56] that BAO and SN observables can be matched to percent-level accuracy between PAge and $\Lambda$CDM.
@@
While the background evolution is matched between PUDF and $\Lambda$CDM, the abundance and equation of state of the unified dark fluid in PUDF are very different from those of dark matter in $\Lambda$CDM. We may expect very different density perturbations of the dark components in the two models. However, density perturbations of the dark components are not directly observable. What can be observed are the density perturbations of baryonic matter and the bending of the light due to gravitational lensing, both of which track the gravitational potential $\phi$ if anisotropic stress can be ignored. The linear growth of $\phi$ in general depends on the total density perturbation $\delta\rho_{\rm tot}$, the total pressure perturbation $\delta p_{\rm tot}$, and the expansion history of the universe [59]. On sub-horizon scales where the gauge-dependence of $\delta\rho_{\rm tot}$ and $\delta p_{\rm tot}$ can be ignored, we may use the Poisson equation to eliminate the dependence on $\delta\rho_{\rm tot}$ [59]. Thus, in models such as PUDF and $\Lambda$CDM where the rest-frame pressure perturbations are assumed to be negligible, the evolution of $\phi$ on sub-horizon scales only depend on the expansion history of the universe. In other words, for background-matched PUDF and $\Lambda$CDM, the linear growth of gravitational potential is also approximately matched. This has been numerically verified in Wang:2024rus [49] where the baryon power spectrum in PUDF was shown to be similar to that in $\Lambda$CDM. In Figure 3 we show that the CMB lensing deflection power spectrum in PUDF and $\Lambda$CDM are similar, too.
@@
Figure 3: Comparison of CMB deflection power spectrum $C_{\ell}^{dd}$ of PUDF and $\Lambda$CDM when the primary-CMB matching condition (2.15) and the $q_{0}$ matching condition (2.17) are applied. The lower panel gives the relative difference($\Delta C_{\ell}^{\mathrm{dd}}=C^{\mathrm{dd}}_{\ell,\mathrm{PUDF}}-C^{\mathrm{dd}}_{\ell,\Lambda\mathrm{CDM}}$).
3 Parameter Inference
@@
In this section we update the parameter inference for PUDF after fixing the minor error in the Boltzmann code. For a fair comparison we use the same combination of CMB + BAO + SN + CC that have been used in Wang:2024rus [49]. These datasets include Planck TTTEEE and lensing likelihoods [60, 61], the Pantheon+ compilation of Type Ia supernovae [62], and the recent Dark Energy Spectroscopic Instrument Data Release 1 (DESI DR1, DESI:2024mwx [8]). More details can be found in Wang:2024rus [49].
@@
The results are listed in Table 2. Compared to the results in Wang:2024rus [49], the updated PUDF parameters all shift towards $\Lambda$CDM parameters. This is because in Wang:2024rus [49] the incorrect usage of HyRec leads to mismatched primary CMB and hence biased cosmological parameters. The updated results still favor $\Lambda$CDM model, but the relative Bayesian evidence derived with the MCEvidence code [63, 64] is updated to $\ln B_{\rm\Lambda CDM,PUDF}=3.75$, much less significant than the previous (incorrect) result $\ln B_{\rm\Lambda CDM,PUDF}=6.27$ in Wang:2024rus [49]. With a Bayesian evidence $\Delta\ln B\approx 3.75$, we may say that $\Lambda$CDM is only mildly preferred by the data and PUDF remains to be an interesting option for future investigation.
@@
data sets $\chi^{2}_{\rm PUDF}-\chi^{2}_{\rm\Lambda CDM}$
CMB+BAO+SN+CC $6.5$
CMB+BAO+CC $0.58$
CMB+SN+CC $1.1$
BBN+BAO+SN+CC $-1.73$
4 Discussion and Conclusions
In this study we show that both the background expansion history and the linear perturbations in the visible sector of the universe can be tuned to be $\Lambda$CDM-like in the PAge-like unified dark fluid model. We derive matching conditions for primary CMB and late-universe observables. Using the primary-CMB matching condition to test the numerical accuracy of the PUDF Boltzmann code, we find a minor error in the code used in Wang:2024rus [49]. After fixing the numerical error we update the parameter inference and find that $\Lambda$CDM is mildly favored by the current CMB+BAO+SN+CC data, with a $\Delta\chi^{2}\approx 6.5$.